A certain mixture is made by mixing two ingredients in the ratio 7 to 3. If it is required to make 40 pounds of mixture, how much of each ingredient will be needed.

Answer by josgarithmetic(39618)

The 40 pounds mixture is understood as 7+3=10 parts. Each part from the contributions of the ratio is then 4 pounds. The "7" ingredient is 7*4=28 parts, and the "3" ingredient is 3*4=12 parts, of the mixture.